OpenStudy (anonymous):

You have a standard number cube. What is the probability of rolling a number less than 3, and then rolling a prime number?

OpenStudy (kropot72):

How many numbers less than 3 are on the cube?

OpenStudy (anonymous):

2

OpenStudy (kropot72):

Yes, there are 2 numbers less than 3 on the cube. So the probability of rolling one of those numbers is 2/6. How many prime numbers are on the cube?

OpenStudy (anonymous):

2 3 5

OpenStudy (kropot72):

Correct. So the probability of rolling a prime number is 3/6. The events are independent, therefore the required probability is the product of the two values of probability: \[\large P=\frac{2}{6}\times\frac{3}{6}=you\ can\ calculate\] Note: The fractional result must be simplified.

OpenStudy (anonymous):

1/6

OpenStudy (kropot72):

That is correct.